OpenStudy (anonymous):

multiply 3y^5/3y^5 x 2z/7x

OpenStudy (stapeleliz):

Is the expression as follows? \[\left(\frac{3y^5}{3t^5}\right)\times\left(\frac{2z}{7x}\right)\] If so, then note that the first fraction reduces to "1".

OpenStudy (anonymous):

yes but 3y^5 on the lower left

OpenStudy (stapeleliz):

(Sorry; that was a typo.)

OpenStudy (anonymous):

no worries